OpenStudy (anonymous):

calculate the sum: 5 Σ 3 i=1

OpenStudy (turingtest):

what is the first term in the series?

OpenStudy (turingtest):

\[\sum_{i=1}^{5}3\]correct?

OpenStudy (anonymous):

yes!!

OpenStudy (turingtest):

what is the first term in the series?

OpenStudy (anonymous):

im not understanding. is it 5?

OpenStudy (anonymous):

i?

OpenStudy (turingtest):

the first term is 3 every term is 3 there are 5 terms \(i\) is not in the summand (that's the part after the sigma), so each term is the same \[\sum_{i=1}^{5}3=5(3)=15\]

OpenStudy (turingtest):

to make it more clear perhaps I should write ti more explicitly\[\sum_{i=1}^{5}=3+3+3+3+3=3(1+1+1+1+1)=3(5)=15\] and in general, for any constant \(k\) we could say that\[\sum_{i=1}^{n}k=kn\]